OpenStudy (anonymous):

Functions f(x) and g(x) are shown below: f(x) g(x) f(x) = 3x2 + 12x + 16 g(x) = 2 sin(2x - π) + 4 Using complete sentences, explain how to find the minimum value for each function and determine which function has the smallest minimum y-value.

OpenStudy (phi):

the minimum of a parabola is at the "vertex" can you find the x value of the vertex ?

OpenStudy (anonymous):

i dont know

OpenStudy (atlas):

find the derivative and equate it to zero............the derivative of a function at minimum and maximum values are zero

OpenStudy (phi):

see http://hotmath.com/hotmath_help/topics/vertex-of-a-parabola.html see the example

OpenStudy (anonymous):

is the vertex 12

OpenStudy (anonymous):

@phi

OpenStudy (phi):

did you find the formula to find the vertex at the site I posted?

OpenStudy (anonymous):

y = ax2 + bx + c.

OpenStudy (phi):

you need to know that. But you also need a formula for the vertex. Look for the sentence right above the Example box.

OpenStudy (anonymous):

y = ax2 + bx + c.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

This site gives it step by step http://www.wikihow.com/Find-the-Vertex-of-a-Quadratic-Equation

OpenStudy (anonymous):

@phi how do find the minim est value of the functiooons.

OpenStudy (phi):

first find the x value of the vertex of the parabola. x = -b / (2a) use that x value in the equation of the parabola to find the y value (this will be the min value you need)

OpenStudy (anonymous):

whats the -b and a value that i plug in @phi

OpenStudy (phi):

you should match these equations \[ f(x) = ax^2 + \ b\ x +\ c \\ f(x)=3x^2 + 12x + 16\] do you see the pattern?

OpenStudy (anonymous):

@phi